### v2

# Complexity in Complexity: Understanding Visual Complexity Through Structure, Color, and Surprise

Abstract
Understanding human perception of visual complexity is crucial in visual cognition. Recently (Shen, et al. 2024) proposed an interpretable segmentation-based model that accurately predicted complexity across various datasets, supporting the idea that complexity can be explained simply. In this work, we investigate the failure of their model to capture structural, color and surprisal contributions to complexity. To this end, we propose Multi-Scale Sobel Gradient which measures spatial intensity variations, Multi-Scale Unique Color which quantifies colorfulness across multiple scales, and surprise scores generated using a Large Language Model. We test our features on existing benchmarks and a novel dataset containing surprising images from Visual Genome. Our experiments demonstrate that modeling complexity accurately is not as simple as previously thought, requiring additional perceptual and semantic factors to address dataset biases. Thus our results offer deeper insights into how humans assess visual complexity.
Keywords: visual complexity; natural images; image segmentation; image gradients; colorfulness; large language models; surprise
11footnotetext: These authors contributed equally to the supervision of the project.
Introduction
Visual complexity is a fundamental attribute of images that reflects the level of detail, intricacy, and variation in visual elements within a scene (?, ?). This perceptual characteristic encompasses multiple dimensions, including the number of lines, density of elements (?, ?), quantity of objects (?, ?), clutter (?, ?), symmetry (?, ?), and variety of colors (?, ?). Understanding visual complexity is crucial across domains, from user interface design (?, ?, ?) to cognitive psychology (?, ?, ?). While deep learning models predict complexity well, their black-box nature (?, ?) limits interpretability, which is essential for applications like education (?, ?, ?) and information visualization (?, ?, ?). Developing interpretable, perception-aligned features remains a key research goal.
A recent study by (?, ?) bridged the gap between handcrafted features and deep learning approaches by proposing a simple two-feature model based on outputs from deep segmentation models (the numbers of segments and semantic classes in the image). Their simple, yet interpretable, model achieved superior performance compared to baseline on many naturalistic and art datasets, supporting the idea that the numbers of segments and classes are generic features that explain complexity well. However, their model fails to take into account the structure and arrangement of components in the image.
Here, we build upon the model from (?, ?), by studying two of its major failure modes. First, we find that datasets with high structural and color regularity require two low-level features to explain complexity: Multi-Scale Sobel Gradient (MSG) and Multi-Scale Unique Color (MUC). MSG captures continuous spatial intensity variations across multiple scales, offering a richer representation of image structure. MUC quantifies color diversity also at multiple scales, and at different color resolutions. The latter feature builds on the concept of “colorfulness” from (?, ?), and provides a more robust and performant measure of chromatic complexity.
We discovered a second failure mode of (?, ?) which concerns the rather underexplored contribution of whole-image, holistic or emergent information to visual complexity. Previous research by (?, ?) demonstrated that participants in a visual complexity experiment exhibited a familiarity bias, perceiving familiar shapes as less complex compared to unfamiliar or novel ones. Similarly, (?, ?) found a negative correlation between visual complexity and familiarity in their experiment. (?, ?) suggests that complexity judgments are context-dependent and that familiarity should be considered when developing a model of visual complexity. Motivated by their findings, we show that holistic, whole-image surprise judgments of scenes contribute significantly to the visual complexity of naturalistic images. We define surprising images as those containing unusual, or contextually novel elements, i.e., the opposite of familiarity. (?, ?, ?). To this end, we introduce a novel dataset, SVG, containing both surprising and randomly sampled images from the well-studied Visual Genome dataset (?, ?) and show that surprise scores generated by a large language model (LLM) explains significant variance in the perceived complexity of images in this dataset.
Methods
Modeling and Evaluation
In this work, we evaluate the effectiveness of sets of features to explain visual complexity on datasets using linear regression models. We follow the procedure in (?, ?), and fit $M$ repetitions of 3-fold cross-validated linear regression. $M$ is determined by the size of the dataset to address statistical variability in smaller datasets. We measure performance using the mean Spearman correlation coefficient across all test splits. We compare the performance of our models to the same baselines in (?, ?): handcrafted features (?, ?, ?) and deep learning approaches (?, ?, ?) for comparison. The code for our analysis, experimental setup, and dataset will be made publicly available upon acceptance.
Datasets
We evaluate our method on four publicly available datasets with human-rated complexity scores: RSIVL (49 images) (Corchs et al., 2016), VISC (800 images) (Kyle-Davidson et al., 2023), Savoias (1400 images across 7 categories) (Saraee et al., 2020), and IC9600 (9600 images across 8 categories) (Feng et al., 2022). We select image subsets to demonstrate the effectiveness of our new features to explain complexity - VISC and IC9600 architecture subset for MSG, Savoias Art and Suprematism subsets for MUC, and Savoias Interior Design and IC9600 Abstract for the combination of MSG and MUC. Finally, we use our novel dataset SVG to show that surprisal contributes significantly to complexity.
SVG: A Dataset of Surprising Images
To the best of our knowledge, there is currently no existing image set with surprising images that can be used to study visual complexity. To fill this gap, we introduce a new dataset containing 100 highly surprising images and 100 (on average, less surprising) images, the latter of which were sampled uniformly at random from the Visual Genome, subject to stratification according to bins of complexity based on IC9600 scores (?, ?). To confirm that our subjectively selected images are more surprising than randomly sampled ones, we compare the distributions of LLM-generated surprise scores between the two subsets. A histogram of these scores shows clear separation: sampled images cluster at lower values, while handpicked images have higher surprise scores (mean: $39.45$ for ordinary, $72.35$ for surprising). A Kolmogorov-Smirnov test confirms a significant difference ($D=0.67,p<0.001$), supporting our hypothesis.
We collect visual complexity ratings from humans using an online experiment on Prolific (?, ?). Participants are shown pairs of images and asked to select the one that is more visually complex, similarly to the procedure in (?, ?). We followed the sampling strategy in (?, ?), and collect $6000=200\times 30$ comparisons by sampling images with probabilities inversely proportional to the total number of times they have already been selected. We ensure that each pairwise comparison is assessed by three different participants. Hence there are in total $18,000$ evaluations. Each participant evaluates $200$ pairwise comparisons in one session. We convert pairwise comparisons into scalar complexity ratings using the Bradley-Terry algorithm (?, ?).
Multi-Scale Sobel Gradient
We propose the Multi-Scale Sobel Gradient (MSG) algorithm (Algorithm 1) to capture structure in images. MSG analyzes spatial intensity variations across multiple resolutions in an RGB image. Note that the Sobel operator is typically applied to grayscale images (?, ?), but we found in ablations that the grayscale version of the algorithm (which first converts the colored images to grayscale) did not perform as well. Hence we use the color version of the algorithm for the rest of this work.
Algorithm 1 Multi-Scale Sobel Gradient
0:  Input image $img\in\mathbb{R}^{H\times W\times 3}$ in RGB format
0:  Scalar MSG score
1:  Normalize image: $img\leftarrow img/255.0$
2:  Define scales $S=\{1,2,4,8\}$ and weights $W=\{0.4,0.3,0.2,0.1\}$
3:  Initialize $MSG\leftarrow 0$
4:  for each scale $s\in S$ and weight $w\in W$ do
5:     Resize image to $\frac{H}{s}\times\frac{W}{s}$: $scaled\leftarrow Resize(img,(\frac{H}{s},\frac{W}{s}))$
6:     for each channel $c\in\{R,G,B\}$ do
7:        $G_{x}=Sobel(scaled[:,:,c],dx=1,dy=0)$
8:        $G_{y}=Sobel(scaled[:,:,c],dx=0,dy=1)$
9:        $mag=\sqrt{G_{x}^{2}+G_{y}^{2}}$
10:        $grad_{c}\leftarrow Mean(mag)$
11:     end for
12:     $s\_grad\leftarrow Mean(grad_{R},grad_{G},grad_{B})$
13:     $MSG\leftarrow MSG+w\cdot s\_grad$
14:  end for
15:  return  $MSG$
Multi-Scale Unique Color
We introduce MUC which intuitively counts the number of unique colors present in an image. We start from the “colorfulness” feature introduced by (?, ?) (which was first applied to image indexing and content querying) and derive MUC by making colorfulness multi-scale. Note, (?, ?) and subsequent works did not provide pseudocode for the colorfulness algorithm - we are the first to do so. MUC has two hyperparameters which control the coarse-graining of spatial resolution and color resolution (bit precision). In the rest of this paper we report results using the best bit precision for each dataset in terms of correlation to complexity (this was typically 7–8 bits). We use a fixed set of spatial scale weights for both MSG and MUC on all datasets.
Algorithm 2 Multi-Scale Unique Color
0:  Image $I$ (RGB), Number of bits to preserve per channel $b$
0:  Multi-Scale unique color score
1:  Define scales $S=\{1,2,4,8\}$ and weights $W=\{0.4,0.3,0.2,0.1\}$
2:  Initialize $MUC\leftarrow 0$
3:  for each scale $s\in S$ and weight $w\in W$ do
4:     Resize image to $\frac{\text{width}}{s}\times\frac{\text{height}}{s}$: $I_{s}\leftarrow\text{resize}(I_{rgb},(\frac{\text{width}}{s},\frac{\text{height}}{s}))$
5:     Calculate bit shift: $shift\leftarrow 8-b$
6:     Quantize image by bit shifting: $I_{q}\leftarrow(I_{s}\gg shift)\ll shift$
7:     Flatten image: $I_{flat}\leftarrow\text{reshape}(I_{q},-1,3)$
8:     Create color index: $idx\leftarrow I_{flat}[:,0]\cdot 2^{16}+I_{flat}[:,1]\cdot 2^{8}+I_{flat}[:,2]$
9:     Count unique colors: $n_{\text{unique}}\leftarrow|\text{unique}(idx)|$
10:     Add weighted contribution: $MUC\leftarrow MUC+(w\cdot n_{\text{unique}})$
11:  end for
12:  return  $MUC$
Generating Surprise Scores using LLMs
In order to have an automated pipeline for predicting complexity, we prompt an LLM using zero-shot Chain-of-Thought (?, ?) to get surprise scores for each image in our SVG dataset. Our prompt is shown in Algorithm 3. We selected gemini-1.5-flash for its ability to provide rapid responses without compromising quality. The model generated surprise scores with step-by-step reasoning, producing structured output in the desired format. We also collect human surprise ratings on the SVG image set to validate our LLM-generated scores using a similar Prolific experiment to that described above.
Algorithm 3 Zero-shot-CoT
Q: Step by step, explain why this image is surprising or not. Consider factors like rare events, or unexpected content. Be precise in your reasoning. Then, on a precise scale from 0 to 100, rate the surprisal of this image.
Provide your reasoning and numeric rating as follows:
Reasoning: [your explanation]
[…]
Results
Table 1: Model performance on 7 datasets and comparison with previous models. The models from previous work are classified as being based on either handcrafted features or Convolutional Neural Networks (CNNs). * for supervised methods indicate their own test set. Bold indicates the best model.
Model/Dataset VISC IC9. Arch. Sav. Art Sav. Suprematism Sav. Int. IC9. Abstract SVG
Handcrafted features
   Corchs 1 (10 features) 0.62 0.66 0.68 0.80 0.85 0.74 0.73
   Kyle-Davidson 1 (2 features) 0.60 0.54 0.55 0.79 0.74 0.69 0.70
CNNs
   Saraee (transfer) 0.58 0.59 0.55 0.72 0.75 0.67 0.72
   Feng (supervised) 0.72 0.92 0.81 0.84 0.89 0.94 0.83
Previous simple model
   $\sqrt{\textit{num\_seg}}+\sqrt{\textit{num\_class}}$ 0.56 0.66 0.73 0.89 0.61 0.66 0.78
Visual features
[…]
Baseline + visual features
   $\sqrt{\textit{num\_seg}}+\sqrt{\textit{num\_class}}+\textit{MSG}$          0.68 $\uparrow 0.13$          0.76 $\uparrow 0.10$ 0.75 0.90 0.79 0.79 0.78
   $\sqrt{\textit{num\_seg}}+\sqrt{\textit{num\_class}}+\textit{MUC}$ 0.62 0.71          0.81 $\uparrow 0.08$          0.94 $\uparrow 0.05$ 0.80 0.76 0.79
   $\sqrt{\textit{num\_seg}}+\sqrt{\textit{num\_class}}+\textit{MSG}+\textit{MUC}$ 0.68 0.77 0.81 0.94          0.87 $\uparrow 0.26$          0.83 $\uparrow 0.17$ 0.79
Baseline + semantic feature
   $\sqrt{\textit{num\_seg}}+\sqrt{\textit{num\_class}}+\textit{Surprise}$ 0.60 0.67 0.74 0.89 0.60 0.67         0.83 $\uparrow 0.05$
Baseline + all features
   $\sqrt{\textit{num\_seg}}+\sqrt{\textit{num\_class}}+\textit{MSG}+\textit{MUC}+\textit{Surprise}$ 0.71 0.78 0.81 0.94 0.87 0.84 0.85
Surprise to the Rescue: Capturing the Complexity Beyond Structural Features
The features we introduced: MSG, MUC, and Surprise, explain additional variance in perceived complexity beyond that captured by segmentation and class counts alone. We present these results in Table 1, showing that our models not only surpass all handcrafted features but also achieves performance comparable to the best supervised model. In the following subsections, we present each feature in turn and describe the datasets on which they improve complexity prediction. We then describe our penultimate model, which combines our three novel features to achieve state-of-the-art or near state-of-the-art performance on all datasets.
Explaining the Failure Modes with a Data-Centric Approach: Colors and Structure
On the VISC dataset, the MSG feature improves the mean Spearman’s rank correlation coefficient from 0.56 to 0.68, surpassing all previous handcrafted features. Similarly, it achieves a notable increase of 0.10 in mean Spearman’s rank correlation coefficient on the IC9. Architecture dataset. These two datasets comprise of mostly real-world natural images of buildings, and indoor and outdoor scenes which are highly textured or regular - for example a grass field, the windows on a highrise, or rows of seats in an airport. A Sobel operator applied to a spatially coarse-grained image can be interpreted as an (a) symmetry detector. Hence MSG detects (a)symmetry and quantifies the regularity of local patches (e.g. windows on a highrise) at multiple spatial scales, making it a well-suited feature for predicting complexity on VISC and IC9. Architecture.
Table 1 shows that MUC improves model performance on Sav. Art dataset, increasing Spearman’s correlation from 0.73 to 0.81, and in Sav. Suprematism, where it improves performance from 0.89 to 0.94. These datasets are comprised of 2D art. The efficacy of MUC in these domains likely stems from its ability to quantify color-based complexity, which may be more crucial for people’s aesthetic judgments than judgments of naturalistic images.
Two datasets, Sav. Interior Design and IC9. Abstract experience very large performance gains (from 0.61 to 0.87 and from 0.66 to 0.83 respectively) when both MSG and MUC features are included in the model. Interior Design is a dataset of synthetic renderings of rooms for IKEA advertisements. The dataset likely benefits from MSG for the same reason as VISC - the images contain regular, highly symmetric elements like books in a bookshelf or panels of a kitchen cabinet. Furthermore, unlike naturalistic images Sav. Interior Design images were designed with color themes in mind for aesthetic reasons, similar to art images which is likely why the Sav. Interior Design dataset benefits from MUC like the Sav. Art dataset. IC9. Abstract on the other hand, consists of real-life photographs. However, the images are taken in an artistic way with aesthetic value in mind, likely increasing the importance of colorfulness in predicting complexity.
Figure 2 presents two examples from the Sav. Interior Design dataset where the baseline model (?, ?) assigns the same complexity due to similar SAM segmentations and FC-CLIP classes, despite differing ground truth complexities. However, the two images differ greatly in their MSG values (20 vs. 92 on a normalized scale of 100), possibly because the lower image has patterns on the bookshelves and blanket which contribute substantially to MSG. The lower (higher) MSG values in the top (bottom) image decrease (increase) the complexity predictions in model, making the predictions much closer to the ground truth complexity ratings.
(a) B: 50, P: 39, G: 40
(b) MSG: 20
(c) B: 50, P: 67, G: 68
(d) MSG: 92
Figure 1: Left column: original images from Sav. Int.. Right column: gradient visualizations. B: baseline prediction using number of segmentations and classes. G = ground truth complexity. P = predicted complexity using baseline and MSG. All values are scaled between 0 and 100. The first image has 177 segmentations and 35 classes, while the second has 185 segmentations and 38 classes. Due to these similarities, the baseline model predicts nearly identical complexity scores for both. However, MSG acts as a latent dimension, refining predictions to better align with ground truth.
Surprise score: 85
Reasoning: The image depicts sheep dressed in jockey silks and participating in a race. This is surprising because it is not a typical activity for sheep. The juxtaposition of animals in a human sport is unexpected and humorous.
Surprise score: 85
Reasoning: The image depicts a skateboarder performing an unexpected maneuver on a beach, resulting in an unusual pose mid-air. The rarity stems from the seemingly uncontrolled backflip and the unusual beach setting for such a stunt.
Figure 2: Racing sheep with toy riders or a flying skater from the SVG dataset illustrate the improvement in complexity predictions when incorporating surprise. The baseline model underestimates the visual complexity, with scores of B: 54 and B: 46, while ground truth values are G: 60 and G: 55. Incorporating surprise scores (85) reduces this gap, yielding adjusted predictions of P: 61 and P: 55, demonstrating the role of surprise in aligning predictions more closely with human perception. Explanations provided by gemini-1.5-flash enhance the interpretability of the assigned surprise scores.
Table 1 shows that including surprise scores generated using gemini-1.5-flash improves Spearman’s correlation on the SVG dataset from 0.78 to 0.83 over (?, ?)’s baseline model. Our model containing surprise scores matches the performance of (?, ?)’s supervised network.
Figure 3: Correlation between residuals (actual complexity - baseline predictions) and surprise scores (SVG).
Furthermore, Figure 3 shows that the surprise scores have a correlation of 0.48 ($p<0.001$) with complexity after regressing out SAM number of segments and FC-CLIP number of objects, showing that surprise captures variance in complexity not explained by the number of segments or objects. These quantitative results are corroborated by participants’ anecdotes on how they rated complexity on the SVG dataset. Many reported that “unusual”, “surprising”, or “weird”, objects or events in the image, as well as judgments that the images are “confusing” or “unreal” caused them to rate the visual complexity higher. Figure 4 presents a selection of responses from participants in our experiment, in which they describe the specific strategies they employed to assess visual complexity. Their responses provide a deeper understanding of the cognitive processes involved in evaluating image complexity and highlight how the element of surprise works as a latent dimension for visual complexity evaluation.
Participant Responses Question: What strategy did you use to rate visual complexity? “Checked if they look confusing or unreal.” “How common the situation was, the angle the image was taken at, the number of colours, the number of subjects and how in focus they were.” “How many different elements there were in each photo and how surprising or unusual the images were.” “The most complex were those with many different elements to them and the most surprising.” “The weird or uncanny images appeared quite complex to me, most of the landscape shots or regular street/traffic scenes didn’t strike me as complex.” “Chose the confusing ones.” “Where there are lots of things to look at. Was there something at a distance in the background? Was the image novel?” “The most complex were the ones with a lot going on in them—so more than one person or something unusual happening. The least complex were the ones of sport or vehicles which just look like generic images.” “If there was a lot going on in the image or if the image was out of the ordinary. Also, images where there were shadows and different types of light.” “I think it was a mix of structure of the image and the elements within it—the more unusual or juxtaposed the mix, the more complex it appeared to me.” “At first, I was thinking about the colours, people, and background. I then thought about how complex the act was in the photo.” “I looked at how busy the picture was, then if it wasn’t too busy I compared them based on what was happening.”
Figure 4: A selection of responses from participants describing the strategies they employed to assess visual complexity. Their responses highlight various factors such as the number of elements, unusual or weird features, level of detail, and how confusing or unreal an image appeared. Words related to surprise, such as ‘unusual,’ ‘confusing,’ and ‘weird,’ have been highlighted in red for emphasis.
Figure 2 shows two examples from the SVG dataset where the model from (?, ?) underpredicts visual complexity. Both images have high (85) surprise scores. Hence our model including surprise scores correctly increases its complexity predictions, making the predictions much closer to the ground truth complexity. Furthermore, one advantage of using LLMs in our automated pipeline for complexity prediction is LLMs offer (often) interpretable, explicit reasoning in natural language (?, ?). Two examples of explanations generated from gemini-1.5-flash using CoT prompting (?, ?, ?), can be seen in Figure 2. In the top image, gemini-1.5-flash successfully identifies the intuitively surprising fact that sheep are racing, an unusual activity for sheep. In the bottom image, gemini-1.5-flash identifies that it’s unusual for the man to be doing an uncontrolled backflip, arguably consistent with human intuition for why the image is surprising.
To explore quantitatively how well our gemini-1.5-flash surprise scores aligned with human judgments of surprise, we computed Spearman’s correlation between gemini-1.5-flash scores and the human surprise scores we collected. We find a high correlation of 0.75. This supports the fact gemini-1.5-flash explanations are good proxies for human judgments of image surprisal which contribute to human perceptions of visual complexity. In summary, gemini-1.5-flash surprise scores improve the predictive power of our complexity model on the SVG dataset and the natural language explanations generated alongside these scores facilitate a clearer understanding of why the image is complex.
The last row of Table 1 shows the performance of our penultimate model including MSG, MUC, surprise as well as the two features from (?, ?). Interestingly, we find combining MSG and MUC with surprise only improves performance on all datasets over using either of these features alone. Furthermore, our penultimate model is significantly more performant than every baseline except (?, ?)’s supervised network which is competitive. Taken together, in contrast to what (?, ?) previously suggested, the complexity of naturalistic images cannot be explained simply using two generic segment and object features. Some datasets require additional features to explain visual complexity fully.
Refer to caption (a) $\sqrt{\textit{num\_seg}}$ = 191 $\sqrt{\textit{num\_class}}$ = 9 MSG = 59 MUC = 922 Surprise = 25 Refer to caption (b) $\sqrt{\textit{num\_seg}}$ = 172 $\sqrt{\textit{num\_class}}$ = 9 MSG = 56 MUC = 820 Surprise = 85 Surprise as the hidden dimension We present a case where both images share similar visual features, including segmentations, classes, MSG, and MUC. The right image, with lower feature values, was initially predicted to have lower complexity. The left image received a complexity score of 60, while the right received 55. However, the element of surprise shifts the outcome, with improved predictions at 52 (72) compared to ground truth values of 25 (74). Despite their similarity, humans may perceive the second image as more complex due to its unusual composition, where multiple traffic signs are clustered together.
Figure 5: Comparison of two images having similar values of visual features but with differing complexity evaluations.
In Figure 5, we present another pair of images where, this time, the addition of surprise scores causes their predicted complexity values to diverge in opposite directions.
Discussion
This work introduces three features to solve the inadequacy of the generic constructs proposed by (?, ?) to explain complexity in certain datasets: multi-scale gradient analysis, multi-scale unique color, and a surprise score. We experimentally show that complexity can be multifaceted - requiring both generic features (segmentation and object counts) and dataset specific features to predict accurately. Our results demonstrate that pixel-level structural and chromatic features, MSG and MUC effectively capture key aspects of visual complexity, with MSG quantifying continuous spatial intensity variations across scales and MUC capturing color diversity, both complementing object-level information from deep segmentation models. Additionally, we introduce surprise, an image-level “cognitive” feature, as a previously underexplored dimension of visual complexity, representing the degree to which the image as a whole deviates from expected objects or events. We introduce a new dataset SVG containing images with unexpected or novel elements to show that surprise contributes significantly to perceived complexity. We constructed our SVG dataset by manually selecting surprising images and hence acknowledge the limitations in scalability and the potential subjectivity of this approach. We hope our preliminary yet confirmatory results will inspire future research to develop larger datasets using objectively defined measure to further validate the link between surprise and complexity.
VISC IC9. Architecture Sav. Art Sav. Suprematism Sav. Int. Design IC9. Abstract
(a) B: 52, P: 62, G: 63
(b) B: 52, P: 44, G: 44
(c) B: 52, P: 66, G: 68
(d) B: 52, P: 43, G: 41
(e) B: 52, P: 92, G: 94
(f) B: 52, P: 38, G: 38
Figure 6: Representative samples from previous datasets show initial complexity predictions ($B=52$), derived from $\sqrt{\textit{num\_seg}}$ and $\sqrt{\textit{num\_class}}$, refined using MSG (VISC, IC9 Architecture), MUC (Sav. Art, Sav. Suprematism), and MSG + MUC (Sav. Int. Design, IC9 Abstract). The enhanced predictions ($P$) align closer to ground truth scores ($G$).
MSG versus Patch Symmetry and Edge Density
Patch symmetry (?, ?, ?, ?) and edge density (?, ?, ?) are two commonly used low-level features for capturing structural information in an image. We run ablation studies to investigate whether MSG is more performant than these two features. We evaluate on all 16 datasets: VISC, RSIVL, SAVOIAS (5 categories - excluding Advertisement and Visualizations due to heavy presence of text), IC9600 (8 categories), and SVG using permutation tests. We compare MSG against patch symmetry because (?, ?) reported that it captures structural information well on the VISC dataset. We compare MSG against Canny edge density because Canny is a standard metric for predicting visual complexity. It has been applied in (?, ?, ?, ?).
MSG performs significantly better than Canny edge density in five datasets. For example, on IC9. Abstract, MSG achieves significantly higher correlation with the complexity compared to edge density (0.517 vs 0.705). Edge density only outperformed MSG on two datasets. Compared to patch symmetry, MSG performs better on 7 datasets. Patch symmetry outperforms MSG on only one dataset. Our ablations show that MSG is currently the most performant structural feature for predicting complexity, outperforming both edge density and patch symmetry across diverse image datasets.
Single versus Multi-Scale
We perform ablation studies to investigate the necessity of multi-scale computations for MSG and MUC. We measure the correlation between complexity and our visual features using both single and multiple scales. We find that multi-scale MSG outperforms single-scale MSG on 6 datasets, while single-scale MSG outperforms multi-scale MSG on just two datasets out of the 16 datasets tested. The advantage of the multi-scale approach was particularly pronounced in Sav. Suprematism, where correlation increased substantially from 0.474 to 0.616. Similarly, multi-scale MUC outperformed single-scale MUC on 5 datasets, while single-scale MUC outperformed multi-scale MUC on only one dataset. The multi-scale versions of MSG and MUC maximize the correlation with the perceived visual complexity.
Memorability and Aesthetics
The existing literature on visual complexity has largely concentrated on identifying low-level image features that define complexity, rather than examining its perceptual implications for the observer. However, as noted by (?, ?), complexity inherently relates to the difficulty of cognitive processing, suggesting that surprising images should be perceived as more complex. We provide first evidence in support of this hypothesis, showing that surprise, as determined by a LLM, contributes significantly to visual complexity on our SVG dataset.
Prior work links image memorability, a closely related concept to complexity (?, ?), to distinctiveness, finding that unique or atypical features enhance retention. (?, ?, ?) find that distinct facial features are more memorable. (?, ?) find that images that deviate from their neighbors in a CNN embedding space have increased recall. We observe a weak positive correlation between human surprisal ratings and both AMNet-predicted ($\rho=0.25,p<0.05$) (?, ?) and ResMem-predicted memorability ($\rho=0.25,p<0.05$) (?, ?), suggesting that surprise may contribute to, or even be a common cause for both judgments of complexity and memorability. Future works should elucidate the relationship between complexity, memorability and surprise.
The relationship between complexity and aesthetics has been extensively studied, though its precise functional form remains uncertain. (?, ?, ?) provides evidence for an inverted U-shape between visual complexity and aesthetics. (?, ?) confirms the ascending part of Berlyne’s inverted-U curve and employs complexity features to predict beauty judgments in images. However, their measure relies solely on visual features like edges and symmetry, while the authors acknowledge that semantic factors, such as familiarity, may also shape perceived visual complexity. In this work, we find preliminary evidence that surprise contributes to complexity. Given the close relationship between complexity and aesthetics, it would be interesting in future works to investigate the extent to which surprise also influences aesthetic, potentially by mediating complexity. Our findings emphasize the multifaceted nature of complexity, requiring varied features across datasets. An incomplete representation of visual complexity may obscure its link to aesthetics, while a more comprehensive approach integrating visual and cognitive-semantic features could enhance understanding.
Future work
While our work uses whole-image, holistic surprise scores, gemini-1.5-flash identifies specific object relationships and object-contexts that make each image more or less surprising. Therefore, it would be interesting for future works to derive surprise scores (or even other semantic features) from object-centric representations such as scene graphs. Our dataset, SVG enables this type of investigation by including human-rated complexity scores, and inheriting region descriptions, and scene graphs from the Visual Genome dataset. These rich annotations open a promising path for developing and validating features that reflect semantic rather than purely visual contributions to complexity.